# Planet-Planet Scattering and ZLK Migration – The Dynamical History of HAT-P-11

### v1

@@ 4 Conclusions
We have proposed a two-step process of planet-planet scattering followed by subsequent ZLK migration to explain the unusual architecture of the HAT-P-11 system. A violent scattering history in which two former planets are ejected from the system is a viable explanation for the high eccentricity and stellar obliquity of HAT-P-11 c. Then, the ZLK mechanism follows naturally with large mutual inclination between HAT-P-11 b and c. We find that traditional ZLK migration with fixed planetary radius cannot reproduce the orbit of HAT-P-11 b, but accounting for thermally-driven radius inflation can. Using N-body simulations, we have verified that this scenario is consistent with all observational constraints as well as the estimated age of the system. The HAT-P-11 system is thus an excellent case study of planet-planet scattering and ZLK migration, both of which are common effects believed to sculpt the architectures of many planetary systems.
@@
HAT-P-11 b adds to a growing census of planets on polar orbits around their host stars (Albrecht et al., 2021) – the exact nature of this peak in the stellar obliquity distribution is not fully understood on a statistical level (Siegel et al., 2023; Dong & Foreman-Mackey, 2023). ZLK migration triggered by the presence of an external planetary perturber has been shown to be theoretically capable of generating the peak of perpendicular planets (Petrovich & Tremaine, 2016), and Vick et al. (2023) demonstrated that if the inner planet starts on an initially misaligned orbit, then perpendicular planets are preferentially produced. They attributed this primordial misalignment to an inclined binary companion torquing the protoplanetary disk (Spalding & Batygin, 2014; Zanazzi & Lai, 2018). We show that planet-planet scattering can lead to similar ZLK initial conditions for single star systems, and hence we demonstrate that the interplay between planet-planet scattering and ZLK migration can easily produce perpendicular planets as well. While stellar obliquity measurements of close-in planets are relatively common, the same cannot be said of the longer-period giants which would constitute said external perturbers (Rice et al., 2021). The possibility of more stellar obliquity measurements of these longer-period planets in the near future will allow more comprehensive analysis of systems characterized by planet-triggered ZLK migration.
@@
Our study demonstrates the importance of considering the evolution of the physical characteristics of a planet during ZLK migration, in addition to the orbital parameters. In the high-eccentricity epochs of ZLK migration, the planet’s envelope can inflate due to tidal heating in the interior. This is highly significant, affecting not only the timescale of ZLK migration but also the equilibrium semimajor axis the planet settles at. This is particularly important in the close companion ZLK cases, since the strength of the ZLK perturbations is relatively much stronger. We show that radius inflation unlocks areas of parameter space which otherwise would be impossible to access through standard ZLK migration with fixed radius. HAT-P-11 is one of the first candidates with two confirmed planets in which ZLK migration is hypothesized to have occurred. A few other systems for which planet-planet ZLK may be possible have begun to emerge (Beust et al., 2012; Petrovich et al., 2018; Bardalez Gagliuffi et al., 2021), and as more such systems are discovered it will be crucial to consider the evolution of the physical parameters of the planet.
@@
While our study highlights the importance of considering physical evolution of the planet, our prescription is a first-order approximation. Given the significant uncertainties in quantities such as internal composition and tidal quality factor present in all exoplanetary systems we have not explored more sophisticated models, and argue that our first-order approximation provides sufficient qualitative insight. However, future work could more deeply consider avenues such as 1) a more sophisticated tidal model – calculating tidal heating and orbital evolution based on dynamical tides, which are more accurate for high eccentricity orbits, 2) coupling tidal heating and radius inflation during orbital evolution (e.g., implementing a lag time such that the radius does not respond instantaneously to tidal forcing, and considering the change in planetary structure when calculating tidal effects (Ogilvie, 2009)) and 3) considering the effects of atmospheric mass loss (Vissapragada et al., 2022).
@@
Finally, we note that while it may be tempting to place constraints on quantities such as the tidal quality factor with this work, this must be done with caution. The final state of HAT-P-11 b depends on the magnitude of tidal forcing compared to the magnitude of the ZLK effect. In practice, this means that it depends on a complex interplay of the initial semimajor axis, tidal quality factor, tidal love number, initial mutual inclination, and internal composition, none of which are well-constrained. To explore such a large parameter space would be a herculean task, and we again stress that this study does not claim to have done so. Rather, our work should be viewed more qualitatively as a proof-of-concept that thermally-driven radius inflation is highly significant towards the final products of planet-planet ZLK migration.
We thank Doug Lin, Yubo Su, Yasmeen Asali, Chloe Neufeld and Yurou (Nina) Liu for insightful discussions.
@@ Appendix A Self-Consistent Spin, Tidal and Dynamical Equations of Motion
with an analagous expression for the quadrupole distortion of body 2 due to body 1. We also consider the acceleration due to the tidal damping of body 1:

### v2

@@ 4 Conclusions
We have proposed a two-step process of planet-planet scattering followed by subsequent ZLK migration to explain the unusual architecture of the HAT-P-11 system. A violent scattering history in which two former planets are ejected from the system is a viable explanation for the high eccentricity and spin-orbit misalignment of HAT-P-11 c. Then, the ZLK mechanism follows naturally with large mutual inclination between HAT-P-11 b and c. We find that traditional ZLK migration with fixed planetary radius cannot reproduce the orbit of HAT-P-11 b, but accounting for thermally-driven radius inflation can. Using N-body simulations, we have verified that this scenario is consistent with all observational constraints as well as the estimated age of the system. The HAT-P-11 system is thus an excellent case study of planet-planet scattering and ZLK migration, both of which are common effects believed to sculpt the architectures of many planetary systems.
@@
HAT-P-11 b adds to a growing census of planets on polar orbits around their host stars (Albrecht et al., 2021) – the exact significance of this peak in the stellar obliquity distribution is sensitive to detection biases and thus has not been definitively confirmed (Siegel et al., 2023; Dong & Foreman-Mackey, 2023). ZLK migration triggered by the presence of an external planetary perturber has been shown to be theoretically capable of generating the peak of perpendicular planets (Petrovich & Tremaine, 2016), and Vick et al. (2023) demonstrated that if the inner planet starts on an initially misaligned orbit, then perpendicular planets are preferentially produced. They attributed this primordial misalignment to an inclined binary companion torquing the protoplanetary disk (Spalding & Batygin, 2014; Zanazzi & Lai, 2018; Gerbig et al., 2024). We show that planet-planet scattering can lead to similar ZLK initial conditions for single star systems, and hence we demonstrate that the interplay between planet-planet scattering and ZLK migration can easily produce perpendicular planets as well. While spin-orbit misalignment measurements of close-in planets are relatively common, the same cannot be said of the longer-period giants which would constitute said external perturbers (Rice et al., 2021). The possibility of more spin-orbit misalignment measurements of these longer-period planets in the near future will allow more comprehensive analysis of systems characterized by planet-triggered ZLK migration.
@@
Our study demonstrates the importance of considering the evolution of the physical characteristics of a planet during ZLK migration, in addition to the orbital parameters. In the high-eccentricity epochs of ZLK migration, the planet’s envelope can inflate due to tidal heating in the interior. This is highly significant, affecting not only the timescale of ZLK migration but also the equilibrium semimajor axis the planet settles at. This is particularly important in the close companion ZLK cases, since the strength of the ZLK perturbations is relatively much stronger. We show that radius inflation unlocks areas of parameter space which otherwise would be impossible to access through standard ZLK migration with fixed radius. HAT-P-11 is one of the first candidates with two confirmed planets in which ZLK migration is hypothesized to have occurred. A few other systems for which planet-planet ZLK may be possible have begun to emerge (Beust et al., 2012; Petrovich et al., 2018; Bardalez Gagliuffi et al., 2021), and as more such systems are discovered it will be crucial to consider the evolution of the physical parameters of the planet. During the preparation of this manuscript Yu & Dai (2024) performed a similar analysis on the WASP-107 system. Their study used a secular orbital evolution code and also accounted for physical processes such as tidal disruption. This work represents another step forward into more nuanced consideration of coupled structure and orbital evolution.
@@
While our study highlights the importance of considering physical evolution of the planet, our prescription is a first-order approximation. Given the significant uncertainties in quantities such as internal composition and tidal quality factor present in all exoplanetary systems we have not explored more sophisticated models, and assert that our first-order approximation provides sufficient qualitative insight. However, future work could more deeply consider avenues such as 1) a more sophisticated tidal model – calculating tidal heating and orbital evolution based on dynamical tides, which are more accurate for high eccentricity orbits, 2) coupling tidal heating and radius inflation during orbital evolution (e.g., implementing a lag time such that the radius does not respond instantaneously to tidal forcing, and considering the change in planetary structure when calculating tidal effects (Ogilvie, 2009)) and 3) considering the effects of atmospheric mass loss (Vissapragada et al., 2022).
@@
Finally, we note that while it may be tempting to place constraints on quantities such as the tidal quality factor with this work, this must be done with caution. The final state of HAT-P-11 b depends on the strength of tidal forcing compared to the magnitude of the precession induced by the ZLK effect. In practice, this means that it depends on a complex interplay of the initial semimajor axis, tidal quality factor, tidal love number, initial mutual inclination, and internal composition, none of which are well-constrained. To explore such a large parameter space would be a herculean task, and we again stress that this study does not claim to have done so. Rather, our work should be viewed more qualitatively as a proof-of-concept that thermally-driven radius inflation is highly significant towards the final products of planet-planet ZLK migration.
We thank the anonymous referee for comments which greatly improved the manuscript. We thank Doug Lin, Yasmeen Asali, Daniel Fabrycky, Konstantin Gerbig, Evgeni Grishen, Yurou (Nina) Liu, Rosemary Mardling, Michael Poon, Chloe Neufeld, and Yubo Su for insightful discussions. GL is grateful for the support by NASA 80NSSC20K0641 and 80NSSC20K0522. This work has benefited from use of the Grace computing cluster at the Yale Center for Research Computing (YCRC). Some plots in this work have made use of the colormaps in the CMasher package (van der Velden, 2020).
Data Availability
The REBOUND C code used to generate all simulations in this work can be found at https://github.com/tigerchenlu98/reboundx/ tree/hatp11_dynamics.
@@ Appendix A Self-Consistent Spin, Tidal and Dynamical Equations of Motion
with an analogous expression for the quadrupole distortion of body 2 due to body 1. We also consider the acceleration due to the tidal damping of body 1: